OpenStudy (anonymous):

PLEASE HELP sqrt - 9 / (3-2i)+(1+5i)

jagr2713 (jagr2713):

\[\huge \color {Blue}{Welcome \to OpenStudy!} \Huge \color{blue}{\star^{ \star^{\star:)}}}\Huge \color{red}{\star^{ \star^{\star:)}}} \]

OpenStudy (anonymous):

choices a . 12 - 9i / 25 B. 12-9i / 7 C. 9+12i / 25 D. 9+12i / 7

OpenStudy (anonymous):

your help is very much appreciated

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (freckles):

So you are doing: \[\frac{\sqrt{-9}}{3-2i}+(1+5i)\]

OpenStudy (anonymous):

the bottom is set up as (3-2i)+(1+5i) so idk if thats the same as that or not

OpenStudy (anonymous):

That looks right though

OpenStudy (freckles):

So you mean the problem is: \[\frac{\sqrt{-9}}{(3-2i)+(1+5i)}\]

OpenStudy (anonymous):

Yes!

OpenStudy (freckles):

no this expression and the expression you gave aren't equivalent it is like saying 1/1+2 is the same as 1/(1+2) and we know 1+2 isn't the same as 1/3 because 1+2 is 3 and 1/3 doesn't equal 3 --- but any who... do you know the square root of 9?

OpenStudy (anonymous):

3

OpenStudy (freckles):

\[\sqrt{-1}=i\] so I'm going to rewrite this as : \[\frac{\sqrt{9} i}{(3-2i)+(1+5i)}=\frac{3i}{(3-2i)+(1+5i)}\]

OpenStudy (freckles):

so on bottom you have 3+1 and -2i+5i

OpenStudy (freckles):

do you know how to find those sums

OpenStudy (anonymous):

No :(

OpenStudy (freckles):

oh I don't believe that 3+1 equals...  how many circles do you see in all